OpenStudy (anonymous):

Simplify the expression. (–2 – 2i)(–4 + 6i) >8 – 4i >8 – 12i >20 – 4i >–4 – 4i

OpenStudy (anonymous):

Do you know the value for i (iota) ?

OpenStudy (anonymous):

just multiply everything and you should get the answer (i*i=-1)

OpenStudy (unklerhaukus):

\[\quad(–2 – 2i)(–4 + 6i) \] \[=-2(–4 + 6i) -2i(-4+6i)\]

OpenStudy (anonymous):

\[\large{(-2-2i)(-4+6i)}\] \[\large{-2(-4) -2i(-4) -2(6i) -2i(6i)}\] \[\large{8 + 8 i - 12 i -12i^2}\]

OpenStudy (anonymous):

>–4 – 4i

OpenStudy (anonymous):

Now since : \[\large{ i = \sqrt{-1} }\] Therefore, \[\large{i^2 = (\sqrt{-1})^2 = -1 }\]

OpenStudy (anonymous):

Solving this we get : 8 - 4i - 12 = -4-4i

OpenStudy (anonymous):

Therefore the answer is : >-4-4i

OpenStudy (anonymous):

thank you all so much, i've got it now! @SheldonEinstein @surdawi

OpenStudy (anonymous):

Nice!

OpenStudy (unklerhaukus):

mistake !

OpenStudy (anonymous):

20 -4i

OpenStudy (anonymous):

Sorry it will be : \[\large{8 + 12 - 4i = 20 -4i}\]

OpenStudy (anonymous):

I am REALLY VERY VERY SORRY for that... Thanks @UnkleRhaukus for pointing it out.

OpenStudy (unklerhaukus):

\[-12i^2=+12\]

OpenStudy (anonymous):

Yeah